40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Wed Jan 25 15:41:29 2017, Lydia, Update, IMC, 29.5 MHz modulation depth measurement plan
|
Wed Jan 25 17:30:03 2017, Koji, Update, IMC, 29.5 MHz modulation depth measurement plan
|
Wed Jan 25 19:39:07 2017, gautam , Update, IMC, 29.5 MHz modulation depth measurement plan
|
Fri Jan 27 21:25:11 2017, Lydia, Update, IMC, 29.5 MHz modulation depth
|
Sat Jan 28 01:25:51 2017, gautam, Update, IMC, 29.5 MHz modulation depth 
|
Mon Jan 30 19:07:48 2017, gautam, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 01:07:20 2017, Lydia, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 13:46:34 2017, rana, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 14:17:48 2017, gautam, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 22:07:13 2017, gautam, Update, IMC, RF AM stabilization box revamp
|
Tue Jan 31 22:28:39 2017, Lydia, Update, IMC, RF AM stabilization box pulled out
|
Wed Feb 1 11:51:19 2017, Koji, Update, IMC, RF AM stabilization box pulled out
|
Wed Feb 1 16:45:56 2017, Lydia, Update, IMC, RF AM stabilizer box Modification Plan
|
Wed Feb 1 20:49:34 2017, rana, Update, IMC, RF AM stabilizer box Modification Plan
|
Wed Feb 1 23:13:30 2017, Lydia, Update, IMC, RF AM stabilizer box Modification Plan
|
Sun Feb 5 21:56:50 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Tue Feb 7 12:01:10 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Tue Feb 7 17:00:55 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 02:14:10 2017, gautam, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 11:41:43 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 13:24:28 2017, rana, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 18:21:21 2017, gautam, Update, IMC, IMC Demod board
|
Fri Feb 10 19:32:15 2017, Koji, Update, IMC, IMC Demod board
|
Sun Feb 12 01:16:57 2017, gautam, Update, IMC, IMC length loop - summary of changes    
|
Mon Feb 13 11:55:14 2017, rana, Update, IMC, IMC length loop - summary of changes
|
Mon Feb 13 13:34:44 2017, gautam, Update, IMC, IMC length loop - bad SMA cable replaced
|
Mon Feb 13 19:44:55 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Wed Mar 1 21:15:40 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Wed Mar 1 23:56:09 2017, gautam, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Thu Mar 2 20:32:18 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Wed Feb 15 23:54:13 2017, gautam, Update, IMC, IMC saga continues... 
|
Fri Feb 17 20:10:18 2017, gautam, Update, IMC, WFS servos turned back on
|
Sat Feb 18 14:09:06 2017, rana, Update, IMC, WFS servos turned back on
|
Sat Feb 18 21:50:48 2017, gautam, Update, IMC, WFS servos turned back on
|
Sun Mar 5 12:41:23 2017, gautam, Update, IMC, WFS servo-steppin  
|
|
Message ID: 12867
Entry time: Sun Mar 5 12:41:23 2017
In reply to: 12840
|
Author: |
gautam |
Type: |
Update |
Category: |
IMC |
Subject: |
WFS servo-steppin |
|
|
I've been sitting on some data for a while now which I finally got around to plotting. Here is a quick summary:
Attachment #1: I applied a step input to the offset of each of the six WFS loops and observed the step response. The 1/e time constant for all 4 WFS loops is <10s suggesting a bandwidth a little above 0.1Hz. However, the MC2 P and Y loops have a much longer time contant of ~150s. Moreover, it looks like the DC centering of the spot on the QPD isn't great - the upper two quadrants (as per the MEDM screen) have ~3x the cts of the lower pair.
I did not (yet) try increasing the gain of this loop to see if this could be mitigated. I accidentally saved this as a png, I will put up the pdf plot
Attachment #2: This is a comparison of the WFS error signals with the loops engaged (solid lines) vs disabled (dashed lines). Though these measurements were taken at slightly different times, they are consistent with the WFS loop bandwidths being ~0.1Hz.
Attachment #3: Comparison of the spectra of the testpoint channels and their DQ counterparts at the same time which are sampled at 512Hz. It does not look like there is any dramatic aliasing going on, although it is hard to tell what exactly is the order of the digital AA filter implemented by the RCG. Further investigation remains to be done... For reference, here are some notes: T1600059, T1400719
GV 7 March 2017 6pm: It looks like we use RCG v2.9.6, so it should be the latter document that is applicable. I've been going through some directories to try and find the actual C-code where the filter coeffs are defined, but have been unsuccessful so far...
Quote: |
I will update with the in-loop error signal spectra, which should give us some idea of the loop bandwidth.
I will look into lowering the sampling rate, and how much out-of-band power is aliasing into the 0-256 Hz band and update with my findings.
|
|
|
|
|